# Takin-VC: Zero-shot Voice Conversion via Jointly Hybrid Content and Memory-Augmented Context-Aware Timbre Modeling

### v1

@@ Abstract
Zero-shot voice conversion (VC) aims to transform the source speaker timbre into an arbitrary unseen one without altering the original speech content. While recent advancements in zero-shot VC methods have shown remarkable progress, there still remains considerable potential for improvement in terms of improving speaker similarity and speech naturalness. In this paper, we propose Takin-VC, a novel zero-shot VC framework based on jointly hybrid content and memory-augmented context-aware timbre modeling to tackle this challenge. Specifically, an effective hybrid content encoder, guided by neural codec training, that leverages quantized features from pre-trained WavLM and HybridFormer is first presented to extract the linguistic content of the source speech. Subsequently, we introduce an advanced cross-attention-based context-aware timbre modeling approach that learns the fine-grained, semantically associated target timbre features. To further enhance both speaker similarity and real-time performance, we utilize a conditional flow matching model to reconstruct the Mel-spectrogram of the source speech. Additionally, we advocate an efficient memory-augmented module designed to generate high-quality conditional target inputs for the flow matching process, thereby improving the overall performance of the proposed system. Experimental results demonstrate that the proposed Takin-VC method surpasses state-of-the-art zero-shot VC systems, delivering superior performance in terms of both speech naturalness and speaker similarity.
@@ 1 Introduction
Zero-shot voice conversion (VC) refers to the task of modifying the timbre of a source speech to match that of a previously unseen speaker, while preserving the original phonetic or linguistic content. This technology has found broad applications in various practical domains [1, 2, 3].
@@
In recent times, zero-shot VC has witnessed great progressions, with numerous state-of-the-art (SOTA) approaches [4, 5, 6, 7, 8, 9] exhibiting impressive results in converting natural and realistic utterances. The key idea behind these methods is to factorize speech into distinct elements, such as linguistic content and timbre elements, and then leverage the source speech content alongside the target speaker timbre to synthesize the desired target speech. In this paradigm, the quality of content and timbre representations, as well as the quality of their disentanglement, significantly impact their final performance. Consequently, numerous studies have sought to improve VC performance by designing more advanced modules [10, 11, 12, 13, 14, 15], information disentanglement approaches [16, 12, 17, 18] and so forth. However, achieving high-quality decoupling of utterances into distinct components remains a challenging task [19, 20, 21, 22], and existing zero-shot VC systems still exhibit subpar performance when handling unseen speakers primarily due to the underlying issues. First, current methods cannot fully mitigate the influence of source speaker timbre during the extraction of linguistic content features, a problem commonly referred to as ”timbre leakage.” Second, they normally employ pre-trained speaker-verification (SV) models to capture target timbre features and cast them as globally time-invariant representations. However, as highlighted in [23], the timbre representations may vary with the linguistic content, rendering the performance of these approaches less optimal. Recently, the advances in large-scale speech language models [24, 25] have tried to tackle this issue by leveraging robust in-context learning capabilities for predicting target speech from concise utterances as prompts. Nevertheless, these methods may suffer from stability issues and error accumulation due to their auto-regressive nature, which can gradually degrade conversion quality.
@@
To address the aforementioned limitations, we introduce Takin-VC, an effective VC framework with advanced modeling of content, timbre and audio quality in a zero-shot fashion. Specifically, we propose a hybrid content encoder guided by neural codec training that integrates the phonetic posterior-grams (PPGs) features and quantized self-supervised learning (SSL) representations from two pre-trained models, i.e., HybridFormer [26] and WavLM [27], so as to precisely capture the linguistic content. For speaker timbre modeling, we first propose a content-aware timbre modeling method that employs cross-attention (CA) to leverage the target voiceprint (VP) features extracted from a pre-trained speaker verification (SV) model [28], with the captured source content. This integration enables our proposed approach to learn target timbre representations associated with source content. Additionally, to further enhance speaker similarity, we advocate a memory-augmented module capable of generating high-quality conditional target inputs for a conditional flow matching (CFM) model [29], ultimately culminating in the synthesis of the target speech using a pre-trained vocoder [30].
@@
To evaluate the performance of the Takin-VC system, we conduct extensive experiments on the both large-scale 500k-hour multilingual (Mandarin and English) and publicly available LibriTTS [31] datasets. Experimental results demonstrate that Takin-VC consistently outperforms state-of-the-art (SOTA) zero-shot VC methods in terms of both speaker similarity and speech naturalness. Notably, Takin-VC achieves significant improvements in both subjective and objective metrics compared with all baseline systems, further validating its effectiveness. For more detailed speech samples, please visit our demo page  [^3] https://everest-ai.github.io/takinvc/.
In summary, the main contributions of this work are outlined as follows:
  • We present Takin-VC, a robust and effective zero-shot VC framework that integrates advanced modeling capabilities for content, timbre, and speech quality. Takin-VC is capable of generating semantically coherent target timbre representations for unseen speakers, resulting in improved speaker similarity and enhanced naturalness/intelligibility.
  • We introduce a hybrid linguistic content encoder that leverages the PPGs and quantized SSL features from the pre-trained HybridFormer and WavLM, with the guidance of neural codec-based training.
  • We propose a context-aware timbre modeling approach based on CA to effectively integrate the source content and target timbre features, bridging the speaker similarity gap between the target speech and ground truth recording.
  • We advocate a memory-augmented module to generate high-quality conditional target inputs for the CFM model, further boosting the speaker similarity performance of our proposed method.
@@ 2 Background
Zero-shot Voice Conversion.
In contrast to previous few-shot [32, 33] and one-shot [12, 34] VC approaches, zero-shot VC presents a more challenging task, as it requires the model to generalize to unseen speakers without any additional training or fine-tuning. In recent years, advancements in deep learning techniques, such as SSL speech models and diffusion models, have led to significant progress in zero-shot VC. SEF-VC [8] learns speaker timbre from reference speech using a CA mechanism and reconstructs waveforms from HuBERT [35] tokens. [6] introduced Diff-HierVC, a diffusion-based hierarchical VC method that uses XLS-R [36] for content extraction and employs two diffusion models to generate high-fidelity converted pitch and Mel-spectrograms. The utilization of robust SSL features, which are rich in phonetic and paralinguistic nuances, has led to improved performance in these methods compared to prior works [37, 38]. Despite these impressive results, SSL-based zero-shot VC approaches [17, 5, 4] may still encounter the timbre leakage problem, as SSL features do not explicitly disentangle timbre information, while diffusion-based VC methods [39, 40] often struggle with poor real-time performance. Another cutting-edge zero-shot VC paradigm [41, 24, 42] involves decoupling speech into semantic and acoustic tokens using neural codecs [43, 44, 45] and SSL speech models [27, 46], subsequently leveraging language models to generate the converted speech. However, these methods still possess great potential for improvement regarding speaker similarity and naturalness/intelligibility.
Flow Matching-based Generative Models.
Recently, flow matching-based generative models [47, 48, 49] have garnered considerable attention in the realm of generative tasks, particularly in the image generation task [50, 51, 52]. These methods focus on approximating the transport probability path from random noise to the target distribution by estimating the associated vector field. By employing a neural ordinary differential equation (ODE), these models learn the optimal transport trajectory and establish a direct link between noise and target samples, which greatly reduces the required number of sampling steps. In contrast to diffusion-based methods [53, 54, 55], flow matching offers improved training stability and real-time performance.
Influenced by this wave of innovation, the speech processing domain has begun to explore flow matching-based generative systems as well. For instance, SpeechFlow [56] leverages a pre-trained generative model using flow matching and masked conditions on extensive untranscribed speech data, enabling effective adaptation to various downstream tasks like speech enhancement, separation, and so forth. ELaTE [57] is a zero-shot TTS system that generates natural laughter by mimicking voice characteristics from an audio prompt and precisely controlling laughter timing and expression through specific input cues. P-Flow [58] utilizes speech prompts for speaker adaptation, integrating a speech-prompted text encoder that generates speaker-conditional representations with a flow matching generative decoder to achieve high-quality, real-time speech synthesis. Nevertheless, the application of flow matching in zero-shot voice conversion (VC) tasks is still in its developmental phase, indicating the urgent need for a stable and efficient flow matching-based zero-shot VC framework.
@@ 3.1 Overivew
As shown in Fig. 1, our Takin-VC system comprises four components: the hybrid linguistic content encoder, memory-augmented & context-aware timbre modeling approach, and CFM model.
@@
[…]
3.2 Hybrid Linguistic Content Encoder
Current zero-shot VC methods typically rely on pre-trained automatic speech recognition (ASR) [59, 60, 61] methods or SSL speech models to extract linguistic content from the original waveform. Nonetheless, both approaches exhibit their respective limitations: the PPGs lacks certain essential paralinguistic nuances, while the latter does not explicitly disentangle timbre information. Therefore, to capture content representations with higher quality, our Takin-VC combines their merits through a neural codec training guided hybrid linguistic content encoder, as shown in Fig. 2.
Figure 2: Content Encoder of Takin VC.
Formally, given an input source speech $x$, the proposed hybrid content encoder first extracts its corresponding PPG and SSL features, denoted as $x_{ppg}$ and $x_{ssl}$, using pre-trained HybridFormer and WavLM, respectively. For our scenario, the HybridFormer is trained on an in-house multilingual corpus of Mandarin and English, while the sixth-layer output features of WavLM are selected as $x_{ssl}$ for further processing.
However, merely relying on the combination of SSL and PPG features is insufficient to achieve optimal VC performance. To further enhance overall performance and address potential timbre leakage, we incorporate a neural codec-based training approach [45] for end-to-end training of the proposed hybrid content encoder, as depicted in the left part of Fig. 2. Concretely, we regard WavLM as the encoder in our neural codec framework and employ a residual vector quantization-based quantizer like [43] to discretize the SSL features. To effectively leverage the PPGs alongside the quantized SSL features, we introduce a simple yet effective fusion module designed to adaptively combine these elements. This module comprises Conv1D layers and ReLU-based gating mechanisms to integrate the SSL and PPG features. The fusion process is formulated as follows:
$x_{s_{cont}}=\text{ReLU}^{\diamond}\left(\text{Conv1d}\left(\alpha_{ssl}\cdot VQ(x_{ssl})+\alpha_{ppg}\cdot x_{ppg}\right)\right)$ (1)
where $\alpha_{ssl}$ and $\alpha_{ppg}$ are learnable hyperparameters, and VQ denotes the vector quantization operation, while Conv1d and $\text{ReLU}^{\diamond}(*)$ represent the convolution and ReLU operations, respectively.
As a consequence, with the guidance of neural codec training, the quality of the fused hybrid SSL and PPG features can be significantly enhanced, resulting in improved naturalness and intelligibility in Takin-VC.
3.3 Context-aware & Memory-augmented Timbre Modeling
3.3.1 Context-aware Timbre Modeling via Cross-attention
Current mainstream VC methods typically regard speaker timbre as a global time-invariant representation [62, 8]. Nevertheless, recent work [23] has uncovered a close correlation between timbre modeling and content information.
Figure 3: The structure of context aware timbre modeling in Takin VC.
Hence, drawing inspiration from this insight, we propose an innovative context-aware timbre modeling approach based on CA. First, we employ a pre-trained SV model to extract a target speaker’s timbre embedding rather than using a global timbre encoder, and then concatenate it with the shuffled Mel-spectrograms of the target speech, denoted as $x_{t_{timb}}$, to minimize the influence of the target content. Subsequently, to learn semantically correlated timbre features that associate the source content with the timbre of the target speaker, we introduce an effective CA-based module. This module takes source content $x_{s_{cont}}$ as the query and $x_{t_{timb}}$ as both the key and value, consisting of a series of linear projection, multi-head CA, layer normalization, and position-wise feed-forward networks (FFN), as detailed in Fig. 3. Finally, we perform interpolation on the extracted features $x_{s_{c}t_{t}}$ to ensure that their dimensionality corresponds to that of the source Mel-spectrogram, thereby facilitating the subsequent training of the CFM model.
@@ 3.3.2 Memory-augmented Timbre Modeling
Figure 4: The structure of memory-augmented timbre modeling in Takin VC.
Since we use a CFM model to reconstruct the source Mel-spectrograms, obtaining high-quality conditional target inputs is quite essential, as they provide key guidance for training the CFM model. To this end, we design an efficient memory-augmented module that adaptively integrates the Mel-spectrogram and VP features of the reference speech, as outlined in Fig. 4. To be specific, our proposed memory-augmented module initially use a Conv1d layer to project the $x_{ref}$ to a latent feature space. Subsequently, we incorporate multiple SA blocks, each containing several group normalization, multi-head SA, and 1D Conv layers, followed by a Conv1d layer and a LeakyReLU activation layer. This design effectively leverages these features in a stable and learnable manner. At the end of memory module, we compute the average vector of the obtained representations across the time dimension to produce the final output $x_{t_{cond}}$. Finally, $x_{t_{cond}}$ is input into the Memory Fusion Layer (a combination of the Gated Activation Layer and FiLM Layer [63]) within the flow matching network to reconstruct the Mel-spectrogram.

### v2

# Takin-VC: Expressive Zero-Shot Voice Conversion via Adaptive Hybrid Content Encoding and Enhanced Timbre Modeling

@@ Abstract
Expressive zero-shot voice conversion (VC) is a critical and challenging task that aims to transform the source timbre into an arbitrary unseen speaker while preserving the original content and expressive qualities. Despite recent progress in zero-shot VC, there remains considerable potential for improvements in speaker similarity and speech naturalness. Moreover, existing zero-shot VC systems struggle to fully reproduce paralinguistic information in highly expressive speech, such as breathing, crying, and emotional nuances, limiting their practical applicability. To address these issues, we propose Takin-VC, a novel expressive zero-shot VC framework via adaptive hybrid content encoding and memory-augmented context-aware timbre modeling. Specifically, we introduce an innovative hybrid content encoder that incorporates an adaptive fusion module, capable of effectively integrating quantized features of the pre-trained WavLM and HybridFormer in an implicit manner, so as to extract precise linguistic features while enriching paralinguistic elements. For timbre modeling, we propose advanced memory-augmented and context-aware modules to generate high-quality target timbre features and fused representations that seamlessly align source content with target timbre. To enhance real-time performance, we advocate a conditional flow matching model to reconstruct the Mel-spectrogram of the source speech. Experimental results show that our Takin-VC consistently surpasses state-of-the-art VC systems, achieving notable improvements in terms of speech naturalness, speech expressiveness, and speaker similarity, while offering enhanced inference speed.
@@ 1 Introduction
Zero-shot voice conversion (VC) aims to modify the timbre of a source speech to match that of a previously unseen speaker, while maintaining the original phonetic content, has found broad applications in various practical domains [1, 2, 3].
@@
The advancement of deep learning techniques has significantly propelled the development of zero-shot VC, with numerous methods [4, 5, 6, 7, 8] exhibiting impressive results in converting natural and realistic speech. The key idea behind is factorizing speech into distinct elements, such as content and timbre elements, and leveraging the source speech content alongside the target timbre to synthesize the desired output. In this paradigm, the quality of content and timbre features, as well as the quality of their disentanglement, critically influences performance. Consequently, various studies have focused on developing advanced modules [9, 10, 11, 12, 13, 14] and information disentanglement approaches [15, 11, 16, 17] to enhance zero-shot VC. However, achieving high-quality decoupling of utterances into distinct components remains challenging [18, 19, 20, 21], with existing systems still exhibiting subpar performance for unseen speakers. Two main issues are prevalent. First, current methods cannot fully mitigate the impact of source timbre during source content extraction, a problem referred to as ”timbre leakage”. Second, these approaches often use pre-trained speaker-verification (SV) models to capture target timbre features as globally time-invariant representations. Nonetheless, such SV embeddings cannot ensure robust timbre modeling and vary with linguistic content [22, 23] which may diminish their effectiveness.
Recently, the progressions in large-scale speech language models [24, 25] have tried to tackle this issue by leveraging robust in-context learning capabilities for converting target speech from concise utterances as prompts. Nevertheless, these methods suffer from stability issues and error accumulation due to their auto-regressive nature, which can gradually degrade conversion quality. Moreover, current state-of-the-art (SOTA) zero-shot VC systems still struggle to simultaneously transfer the paralinguistic characteristics in highly expressive speech, such as crying, breathing, and emotional nuances, thus limiting their effectiveness and practical applicability.
@@
In this paper, we propose Takin-VC, a novel expressive zero-shot VC framework that delivers advanced modeling of content, timbre, and speech quality in a zero-shot fashion. To be specific, we introduce an adaptive fusion-based hybrid content encoder that seamlessly combines the strengths of phonetic posterior-grams (PPGs) and self-supervised learning (SSL)-based representations derived from pre-trained HybridFormer [26] and WavLM [27]. This integration enables the precise extraction of linguistic content while simultaneously enriching paralinguistic elements. For timbre modeling, we first advocate a memory-augmented module capable of generating high-quality conditional target timbre inputs for our conditional flow matching (CFM) model. To further enhance speaker similarity, a context-aware timbre modeling module based on an efficient cross-attention (CA) mechanism is presented. This module effectively aligns and fuse the extracted source content and target timbre features, rather than solely using the source linguistic content as the conditional input for CFM. Conditioned on these features, the predicted outputs of the CFM model are ultimately fed into a pre-trained vocoder [28] to synthesize the target speech.
@@
Experiments conducted on both large-scale 500k-hour multilingual (Mandarin and English) and small-scale LibriTTS [29] datasets demonstrate that Takin-VC consistently outperforms several SOTA zero-shot VC methods in speech naturalness, expressiveness, speaker similarity, and real-time performance. Notably, Takin-VC achieves significant improvements in both subjective and objective metrics compared to all baseline systems, further validating its effectiveness and robustness. For more detailed speech samples, please visit our demo page  [^3] https://anonymous.4open.science/w/takin-vc-0CD8/. In summary, the primary contributions of this work are as follows:
  • We present Takin-VC, a novel expressive zero-shot VC framework. To the best of our knowledge, this is the first approach capable of simultaneously transforming the source timbre to arbitrary unseen speakers while effectively maintaining the paralinguistic characteristics of highly expressive speech.
  • We introduce an adaptive hybrid content encoder that employs an adaptive feature fusion module to implicitly integrate PPGs and quantized SSL features in a learnable manner, thereby capturing precise linguistic elements with enriched paralinguistic characteristics.
  • We propose memory-augmented and content-aware modules to enhance timbre modeling. The former aims to extract high-quality target timbre conditions, while the latter focuses on generating fused features that align and leverage target timbre embeddings with source content for the conditional flow matching model.
@@ 2 Background
2.1 Zero-shot Voice Conversion
Recent progressions in deep learning techniques, such as SSL-based speech models [30, 27, 31] and diffusion models [32, 33], have greatly advance zero-shot VC. SEF-VC [34] utilizes a CA mechanism to extract timbre features and reconstruct waveforms from HuBERT [30] tokens, while [6] proposes a diffusion-based hierarchical VC method using XLS-R [35] for content extraction and dual diffusion models for generating pitch and Mel-spectrograms. Despite these innovations, SSL-based zero-shot VC methods [16, 5, 4] are likely to encounter the timbre leakage challenge, as SSL features do not explicitly disentangle timbre features. Likewise, diffusion-based approaches [36, 37] suffer from suboptimal real-time performance. Another emerging paradigm [38, 24, 39] involves decoupling speech into semantic and acoustic tokens using neural codecs [40, 41, 42] and SSL-based models, subsequently using language models to generate converted speech. While these approaches mark impressive results, current SOTA VC methods still have considerable room for improvement in achieving better speaker similarity and naturalness. Besides, they continue to face difficulties in faithfully and simultaneously reproducing the paralinguistic characteristics of highly expressive speech.
2.2 Flow Matching-based Generative Models
Flow matching-based generative models [43, 44, 45] have recently emerged as a powerful solution for generative tasks. By estimating vector fields to approximate the transport path from noise to the target distribution, these models employ neural ordinary differential equations (ODEs) to learn optimal transport trajectories. Compared to diffusion-based methods [46, 47], flow matching offers improved training stability and real-time performance by enabling direct noise-to-sample mapping while significantly reducing sampling steps. In the speech processing domain, flow matching-based systems are emerging as a promising paradigm. SpeechFlow [48] uses a pre-trained flow matching model with masked conditions on large-scale untranscribed speech data, facilitating speech enhancement and separation tasks. P-Flow [49] adopts speech prompts for speaker adaptation, integrating a speech-prompted text encoder and a flow matching decoder to enable high-quality and real-time speech synthesis. Despite these advancements, the application of flow matching in zero-shot VC remains nascent, underscoring the need for developing a stable and efficient flow matching-based zero-shot VC framework.
@@ 3.1 Overivew
As shown in Fig. 1, our Takin-VC system primarily comprises three key components: an adaptive hybrid content encoder, a memory-augmented context-aware timbre modeling approach, and a conditional flow matching-based decoder.
@@
In detail, the objective of the adaptive hybrid content encoder is to precisely capture linguistic characteristics enriched with paralinguistic elements, denoted as $X_{s_{cont}}$. To achieve this, an adaptive feature fusion module on top of the hybrid content encoder is presented to effectively leverage the complementary strengths of PPG and quantized SSL representations in a learnable fashion. Regarding timbre modeling, we first propose a memory-augmented module that incorporates a stack of convolution, activation, and self-attention layers to extract high-quality target timbre conditions $X_{t_{tcond}}$ for the CFM model. To further improve timbre modeling capabilities, a cross-attention-based context-aware module is presented to generate fused representations $X_{s_{ct_{t}}}$ that effectively integrate $X_{s_{cont}}$ with target timbre. Finally, to enable stable training and accelerate the reference speed, we design a CFM model that consists of multiple UNet [50] blocks to reconstruct the source Mel-spectrograms conditioned on $X_{s_{c}t_{t}}$ and $X_{t_{tcond}}$, followed by a pretrained Bigvgan vocoder to synthesize the desired target speech.
3.2 Adaptive Hybrid Content Encoder
Current mainstream zero-shot VC systems typically use pretrained automatic speech recognition (ASR) [51, 52, 53] or SSL-based speech models to capture linguistic content from the original waveform. However, they both have inherent limitations: ASR-derived PPGs lack sufficient paralinguistic elements, whereas SSL-based models do not explicitly disentangle timbre information. To address these flaws, we propose an adaptive fusion-based hybrid content encoder within the Takin-VC framework, integrating the merits of both approaches.
Formally, given an input source speech $X$, our adaptive hybrid content encoder separately encodes its corresponding PPG and SSL features, denoted as $X_{p}$ and $X_{s}$, using pre-trained HybridFormer and WavLM, respectively. To alleviate potential timbre leakage, a residual vector quantization (RVQ) based quantizer of EnCodec [40] is applied to discretize $X_{s}$, resulting in $\tilde{X_{s}}$. Additionally, we introduce a gradient-driven adaptive feature fusion module to further reduce timbre leakage and effectively integrate the complementary benefits of PPG and SSL features. Unlike conventional element-wise addition for feature fusion, the proposed strategy first processes the quantized WavLM features through a multi-layer projection module comprising a one-dimensional convolutional (Conv1d) layer followed by a LeakyReLU activation function, with the negative slope empirically set to 0.2. Temporal interpolation is then applied to ensure dimensional alignment with the PPG features, and the resulting WavLM representations are employed as coefficients for element-wise multiplication with the PPGs:
$X_{s_{cont}}=\text{LeakyReLU}(\text{Conv1d}(\tilde{X_{s}}))\cdot X_{p}$ (1)
where Conv1d denotes the 1D convolutional layer. By this means, as gradients propagate back to Formula 1 during training, the limited representation of paralinguistic nuances within the PPG features results in larger gradient magnitudes for these elements. Since the PPGs are fixed before training, the gradients primarily affect the adaptive fusion module associated with the quantized WavLM features. As a consequence, this gradient-driven adjustment dynamically optimizes the weights of the quantized WavLM features in an implicit way, thereby amplifying the representation of paralinguistic elements in the combined feature space, improving overall content modeling capabilities, and significantly reducing the risk of voiceprint leakage.
3.3 Enhanced Timbre Modeling
@@ 3.3.1 Memory-augmented Timbre Modeling
Figure 2: Schematic of the memory-augmented module.
To capture high-quality target timbre conditions for the CFM model, we propose an efficient memory-augmented module that adaptively integrates the shuffled Mel-spectrogram and VP features of the reference speech, as outlined in Fig. 2.
Detailed, we extract the Mel-spectrograms from randomly segmented reference waveforms originating from the same speaker as the source speech. The individual frames of these Mel-spectrograms are then shuffled to preserve essential timbre characteristics while minimizing the influence of the source speech content. Subsequently, a lightweight pre-trained SV model [^4] https://modelscope.cn/models/iic/speech_campplus_sv_zh_en_16k-common_advanced is utilized to extract timbre embeddings from the reference speech. These embeddings are then concatenated with the shuffled Mel-spectrograms, resulting in the target timbre representations, referred to as $X_{t_{timb}}$. To refine these concatenated features, our proposed memory-augmented module that begins by employing a Conv1d layer to project the captured features and then incorporates four SA blocks, each comprising a group normalization layer, multi-head SA mechanism, a Conv1d layer, and a shortcut connection operation. The resulting features are then subjected to a temporal averaging operation, followed by the application of a FiLM layer [54] to perform affine feature-wise transformation, producing the conditional target timbre inputs $X_{t_{tcond}}$.
3.3.2 Context-aware Timbre Modeling
Speaker timbre features have long been viewed as global and time-invariant representations [55, 34, 23]. However, recent studies [22] have revealed a close interdependence between timbre modeling and content information. Hence, drawing inspiration from this insight, we propose an innovative context-aware timbre modeling approach based on advanced cross-attention mechanism.
Figure 3: Schematic of the context-aware module.
As illustrated in Fig. 3, the CA-based module is designed to generate semantically aligned timbre features that harmonize the source linguistic content with the target timbre. Concretely, the CA-based module consists of a series of linear projection layers, multi-head cross-attention layers, layer normalization, and position feed-forward network (FFN), which can effectively facilitate the integration of $X_{s_{cont}}$ and $X_{t_{timb}}$. The source content $X_{s_{cont}}$ is used as the query, while the target timbre $X_{t_{timb}}$ serves as both the key and value. Finally, the extracted features $X_{s_{c}t_{t}}$ are interpolated to ensure dimensional compatibility with the ground truth, i.e., the source Mel-spectrogram, facilitating the subsequent training of the CFM model.